# DeepTrace: Learning to Optimize Contact Tracing in Epidemic Networks with Graph Neural Networks

### v3

@@ V-C DeepTrace for COVID-$19$ Epidemic Networks
In this section, we conduct experiments on COVID-$19$ pandemic data in Taiwan and Hong Kong to evaluate the performance of Algorithm DeepTrace. A summary of these epidemic network datasets is presented in TABLE VII. We refer to $G_{n}$ as the epidemic network in the following, where nodes in $G_{n}$ represent confirmed cases, and each edge in $G_{n}$ implies that there is close contact between two confirmed cases. We compare Algorithm DeepTrace to estimators using (2), and since these two clusters are relatively small, the computation of the ML estimator (2) is feasible.
@@
TABLE VII: A summary of COVID-$19$ epidemic network datasets for Algorithm DeepTrace.
@@ V-C1 Contact Tracing for COVID-$19$ Pandemic in Taiwan
Fig. 9 is the largest contact tracing networks constructed with the collected data with $92$ confirmed cases. Case 22595 is the reported superspreader in this contact tracing network. As Case 22595 is also the earliest discovered case, we consider it as the index case for contact tracing. Using the trained GNN in Algorithm DeepTrace, we applied it to this network and detected that Case 22595, which is colored red, is the most likely superspreader. These findings align with the actual superspreaders reported by the Taiwanese authorities.
@@ V-C2 Contact Tracing for COVID-$19$ Pandemic in Hong Kong
As Hong Kong adhered to a strict ’Zero-COVID’ policy with rigorous quarantine and testing measures, its population remained largely unaffected by the COVID-19 virus until January 2022. The introduction of the Omicron variant to this largely susceptible population led to one of the world’s highest death tolls in February and March 2022, despite Hong Kong’s early and widespread access to vaccines and the lower lethality of Omicron compared to previous variants [46, 47]. This unique situation, not previously observed in other parts of the world, presents a real-world scenario where the SI spreading model can be applied to evaluate the performance of Algorithm DeepTrace using training data from Hong Kong’s Omicron outbreak.
We first consider an epidemic network with four connected components in Hong Kong from 31 December 2021 to 22 January, the data for which is from [47]. This epidemic network contains 102 SARS-CoV-2 Omicron infections associated with four imported cases (nodes 1, 29, 70, and 100) in the connected components, which serve as ground truths of superspreaders. In this network, we use DeepTrace to iteratively identify the most likely superspreader as described by (2) as the forward contact tracing network expands. The illustration of the epidemic network and contact tracing process is depicted in Fig. 10. Initially, we select nodes $12$, $36$, $85$, and $102$ as the index cases and the starting points for the forward contact tracing process. As the contact tracing network grows, the most likely superspreader identified by DeepTrace shifts from node $12$ to node $1$, from node $36$ to node $35$ and finally to node $29$, from node $85$ to node $84$ and finally to node $70$, and from node $102$ to node $100$ in the four connected components, respectively. The final detected superspreaders (nodes $1$, $29$, $70$, and $100$) match the most likely superspreader given by (2) in the four connected components. Additionally, nodes $1$, $29$, and $100$ are also the true superspreaders, while node $70$ is just one hop away from the true superspreader, node $68$. These results demonstrate that DeepTrace performs well in tracing the superspreaders.
@@
Next, we consider utilizing DeepTrace on another epidemic network in Hong Kong where the ground truths of superspreaders are unknown. Following [46], we retrieved data of confirmed Omicron cases in Hong Kong from January 31, 2022, to February 3, 2022, from the Hong Kong government’s public sector open data portal at https://data.gov.hk/en-data/dataset/hk-dh-chpsebcddr-novel-infectious-agent. We used this data to construct epidemic networks, where each node represents a confirmed case and an edge connects two nodes if the cases were known to visit the same building at any time. We used the earliest reported case in [46] as the initial index case of the outbreak to conduct forward contact tracing. Algorithm DeepTrace was then employed iteratively to identify the superspreaders as the contact tracing networks expanded.
@@
Fig. 11 presents the contact tracing networks using the retrieved data, illustrating how the iterative outputs of DeepTrace eventually approaches the most likely superspreader given by (2) as the forward contact tracing network expands. For example, in the largest spreading cluster shown in Fig. 11, case 12611 is considered the index case and the initial superspreader. We initiate the forward contact tracing process from this case. As the contact tracing network grows, the most likely superspreader identified by DeepTrace shifts from case 12611 to case 12747, and finally to case 12825, which is exactly the most likely superspreader given by (2) in the entire epidemic network. Therefore, even without prior knowledge of the true superspreaders, DeepTrace can effectively identify the most likely superspreaders. By focusing treatment or implementing quarantine measures on these identified superspreaders, we may be able to suppress further propagation of the epidemic.
@@
Figure 10: Illustration of contact tracing with DeepTrace for an early Omicron outbreak in Hong Kong from 31 December 2021 to 22 January. The blue nodes represent infected nodes. The nodes in the gray boxes are considered the index cases of the Omicron outbreak. The nodes in the red boxes are the true superspreaders. The nodes in pink are superspreaders detected in the subgraph of the epidemic network during contact tracing. The nodes in red are the most-likely superspreaders in the spreading clusters in the epidemic network. The red arrows indicate the movement of superspreaders detected by DeepTrace as contact tracing continues.
@@
Figure 11: Illustration of contact tracing using Algorithm DeepTrace for an early Omicron outbreak in Hong Kong in 2022. The blue nodes represent infected nodes. The nodes in the boxes are the index cases of the Omicron outbreak. The nodes in pink are superspreaders detected in the subgraph of the epidemic network during contact tracing. The nodes in red are the most-likely superspreaders in the spreading clusters in the epidemic network. The red arrows indicate the movement of superspreaders detected by DeepTrace as contact tracing continues.
@@ VI Conclusion
Digital contact tracing remains an open issue in tackling pandemics. In this paper, we have formulated the backward contact tracing problem as a maximum likelihood estimation of the epidemic source, which is crucial for identifying superspreaders in epidemic outbreaks. Given an index case, the contact tracer can explore the epidemic network using breadth-first search or depth-first search to compute the graph center of the instantaneous contact tracing subgraph, converging to the optimal likelihood estimate for certain graph topologies. We proposed Algorithm DeepTrace, which employs a graph neural network with semi-supervised learning, incorporating pre-training and fine-tuning phases for general and potentially large graphs through message passing. This approach is particularly beneficial for massive graph inference in computational epidemiology and is connected to online graph exploration, whose maze-solving perspective offers promising avenues for contact tracing with low-complexity greedy algorithms. We validated the effectiveness of our methodology by modeling and analyzing real data from COVID-19 superspreading events in Hong Kong and Taiwan, demonstrating that our contact tracing algorithm can outperform existing heuristic methods.
To further enhance digital contact tracing, future work will focus on exploring advanced data-driven online graph exploration techniques and adaptive learning algorithms to handle dynamic and evolving epidemic networks. Additionally, we aim to find efficient decomposition techniques to break down the underlying optimization problem into smaller subproblems suitable for machine learning. Extending Algorithm DeepTrace to incorporate real-time data streams and improve scalability for larger networks will be crucial. Lastly, our approach, which focuses on learning to optimize contact tracing, can be refined and incorporated into scalable surveillance systems, enhancing the monitoring and mitigation of future pandemics.
@@ A-A Proof of Theorem 1
Let $\mathbb{G}_{N}$ be a $d$-regular tree and $\overline{v}\in\mathbb{G}_{N}$ satisfies (5), we assume that the maximum distance from $\overline{v}$ to all leaves nodes is $k+1$. Then by (5), there are only two possible distances from $\overline{v}$ to all leaves, which are $k$ and $k+1$. For simplicity, let $T_{d,k}$ denote the perfect $d$-regular tree with $k$ levels, and we assume that the level of the root is 00. We denote $t^{u}_{v}$ as the size of the subtree rooted at $v$ after removing the edge $\{(u,v)\}$ from $\mathbb{G}_{N}$.
We have the following properties of $\mathbb{G}_{N}$:
@@ A-C Proof of Property 2 of $\mathbb{G}_{N}$ used in Theorem 1
Let $\mathbb{G}_{N}$ be a $d$-regular tree and $\overline{v}\in\mathbb{G}_{N}$ satisfies (5), we assume that the maximum distance from $\overline{v}$ to all leaves nodes is $k+1$. Let $\overline{v}$ be the root of $\mathbb{G}_{N}$, and $v\in\mathbb{G}_{N}$ be an arbitrary node. If $v$ is a leaf or the root of $\mathbb{G}_{N}$, then the subtree rooted at $v$ satisfies (5).
If $v$ is not a leaf or the root of $\mathbb{G}_{N}$, then we prove this property by contradiction. Assume that the subtree rooted at $v$, denoted as $T_{v}^{\overline{v}}$ does not satisfy (5), then there are two leaves $u,w\in T_{v}^{\overline{v}}$, satisfy $|\textup{dist}(u,v)-\textup{dist}(w,v)|>1$. since $\mathbb{G}_{N}$ is a tree, we have $\textup{dist}(\overline{v},u)=\textup{dist}(\overline{v},v)+\textup{dist}(v,u)$ and $\textup{dist}(\overline{v},w)=\textup{dist}(\overline{v},v)+\textup{dist}(v,w)$ which implies $|\textup{dist}(\overline{v},u)-\textup{dist}(\overline{v},w)|>1$ and contradicts to our assumption.
@@ A-D Proof of Theorem 2
Note that if a node in $\mathbb{G}_{N}$ with weight strictly less than $n/2$, then the node is the unique ML estimator on contact tracing network $\mathbb{G}_{N}$ [16, 17, 15]. Hence we have
$j_{1}+1=2d_{1}-1.$
Moreover, since $(v=v^{*}_{j_{2}},v^{*}_{j_{2}+1})\in S_{2}$, we have
$j_{2}+1$ $=2d_{2}-1$
$>2(j_{1}+1)-1$
$=2j_{1}+1.$
@@ A-E Proof of Lemma 1
Since the DFS contact tracing starts from the index case $v^{1}$, we can assume that the index case $v^{1}$ is the root of the DFS tree. Now, we are able to distinguish parent-node and child-node in each pair $(v^{*}_{j},v^{*}_{j+1})\in S$. Due to the property of the DFS tracing, a node is fully visited when all its descendants are fully visited; we have the following observation: If $(v^{*}_{j}=u,v^{*}_{j+1}=v)\in S_{2}\cup S_{3}$ and $v$ is the parent of $u$, then there is no other $i>j$ such that $(v^{*}_{i}=v,v^{*}_{i+1}=u)\in S_{2}\cup S_{3}$. Moreover, we can observe that:
@@
Since the estimated superspreader changes from $u$ to $v$ only when $P(\mathbb{G}_{N}\mid v)>P(\mathbb{G}_{N}\mid u)$. We can conclude that both type two and type three transitions must be followed by a type one transition, which means the size of the type one transition is more than type two and type three. Hence, we have $|S_{1}|\geq|S_{2}|+|S_{3}|$. Moreover, from the above two properties, we have the fact that for each $(u,v)\in S_{3}$, there is a corresponding $(v,u)\in S_{2}$, i.e., $|S_{2}|\geq|S_{3}|$. Combining the results above, we have $|S_{1}|\geq|S_{2}|\geq|S_{3}|$.
Since $|S_{1}|+|S_{2}|+|S_{3}|=N-1$ and $|S_{2}|\geq|S_{3}|$, we can deduce that $|S_{1}|\geq\frac{N-1}{2}$ and $|S_{3}|\leq\frac{N-1}{4}$.
@@ A-F Proof of Theorem 3
As shown in (2), solving MLE problem needs the information of all the permitted permutations for all the nodes in $G_{n}$. Since a newly infected node can only be infected by one of its infected neighbors, $v_{i+1}$ must be the neighbor of $v_{i}$ in any permitted permutation $\sigma$, and thus calculating $p(\sigma\mid v)$ with (4) is actually a recursive process of aggregating the information of degree $d(v_{l})$ and $\Phi_{l}$ from the $l$th layer’s neighbors of node $v$ in $\sigma$, which can be expressed as
@@
Note that (13) demonstrates the process of iterative aggregation to calculate $\mathbb{P}(G_{n}\mid v)$ for node $v$. Since $\mathbb{G}$ is connected, the shortest distance between any two nodes satisfies $\textup{dist}(u,v)\leq\textup{diam}(G)$. Thus, after a number of $\textup{diam}(G)$ steps of aggregation and combination, each node can obtain the feature information of all nodes in the graph. Furthermore, it is evident that both $f(\cdot)$ and $g(\cdot)$ are continuous functions. Consequently, based on the universal approximation theorem, these functions can be approximated to any desired degree of precision using standard multilayer perceptrons with hidden layers and non-constant, monotonically increasing activation functions [42].

### v4

@@ V-C DeepTrace for COVID-$19$ Epidemic Networks
In this section, we conduct experiments on COVID-$19$ pandemic data in Taiwan and Hong Kong to evaluate the performance of Algorithm DeepTrace. A summary of these epidemic network datasets is presented in TABLE V. We refer to $G_{n}$ as the epidemic network in the following, where nodes in $G_{n}$ represent confirmed cases, and each edge in $G_{n}$ implies that there is close contact between two confirmed cases. We compare Algorithm DeepTrace to estimators using (2), and since these two clusters are relatively small, the computation of the ML estimator (2) is feasible.
TABLE V: A summary of COVID-$19$ epidemic network datasets for Algorithm DeepTrace.
@@ V-C1 Contact Tracing for COVID-$19$ Pandemic in Taiwan
Fig. 9 is the largest contact tracing network constructed with the collected data with $92$ confirmed cases. Case 22595 is the reported superspreader in this contact tracing network. As Case 22595 is also the earliest discovered case, we consider it the index case for contact tracing. Using the trained GNN in Algorithm DeepTrace, we applied it to this network and detected that Case 22595, colored red, is the most likely superspreader. These findings align with the actual superspreaders reported by the Taiwanese authorities.
@@ V-C2 Contact Tracing for COVID-$19$ Pandemic in Hong Kong
Due to Hong Kong’s strict “Zero-COVID" policy with rigorous quarantine and testing, its population remained largely unaffected by COVID-19 until January 2022. However, the arrival of the Omicron variant led to one of the world’s highest death tolls in early 2022, despite widespread vaccine access and Omicron’s lower lethality [48, 49]. This unique scenario allows for real-world application of the SI spreading model and evaluation of Algorithm DeepTrace using data from Hong Kong’s Omicron outbreak. We first consider an epidemic network with four connected components in Hong Kong from 31 December 2021 to 22 January, the data for which is from [49]. This epidemic network contains 102 SARS-CoV-2 Omicron infections associated with four imported cases (nodes 1, 29, 70, and 100) in the connected components, which serve as ground truths of superspreaders.
DeepTrace is used to iteratively identify the most likely superspreader, as outlined in (2), while the forward contact tracing network expands. Fig. 10 illustrates the epidemic network and contact tracing process. Initially, nodes $12$, $36$, $85$, and $102$ serve as index cases for the tracing process. As the network grows, DeepTrace identifies the most likely superspreaders, shifting from node $12$ to $1$, from $36$ to $35$, from $85$ to $84$ and then $70$, and from $102$ to $100$ across the four connected components. The final detected superspreaders (nodes $1$, $29$, $70$, and $100$) correspond to the predictions from (2). Nodes $1$, $29$, and $100$ are confirmed superspreaders, while node $70$ is one hop away from the true superspreader, node $68$. These results corroborate the efficiency of DeepTrace.
@@
Next, we consider utilizing DeepTrace on another epidemic network in Hong Kong, where the ground truths of superspreaders are unknown. Following [48], we retrieved data of confirmed Omicron cases in Hong Kong from January 31, 2022, to February 3, 2022, from the Hong Kong government’s public sector open data portal at https://data.gov.hk/en-data/dataset/hk-dh-chpsebcddr-novel-infectious-agent. We constructed epidemic networks using this data, where each node represents a confirmed case and edges connect nodes that visited the same building. The earliest reported case in [48] served as the initial index case for forward contact tracing, with DeepTrace iteratively identifying superspreaders as the network expanded.
@@
Fig. 11 illustrates the contact tracing networks built from the retrieved data, showing how DeepTrace’s iterative outputs converge on the most likely superspreader from (2) as the network expands. For instance, in the largest spreading cluster, case 12611 is the index case and initial superspreader. As the tracing network develops, DeepTrace identifies the superspreader, shifting from case 12611 to 12747 and finally to case 12825, which matches the prediction from (2). This demonstrates that DeepTrace effectively identifies likely superspreaders without prior knowledge of the true ones. Targeting these identified superspreaders with treatment or quarantine measures could help suppress further epidemic spread.
@@
Figure 10: Illustration of contact tracing with DeepTrace for an early Omicron outbreak in Hong Kong from 31 December 2021 to 22 January. The blue nodes represent infected nodes. The nodes in the gray boxes are considered the index cases of the Omicron outbreak. The nodes in the red boxes are the true superspreaders. The nodes in pink are superspreaders detected in the subgraph of the epidemic network during contact tracing. The nodes in red are the most likely superspreaders in the spreading clusters in the epidemic network. The red arrows indicate the movement of superspreaders detected by DeepTrace as contact tracing continues.
@@
Figure 11: Illustration of contact tracing using Algorithm DeepTrace for an early Omicron outbreak in Hong Kong in 2022. The blue nodes represent infected nodes. The nodes in the boxes are the index cases of the Omicron outbreak. The nodes in pink are superspreaders detected in the subgraph of the epidemic network during contact tracing. The nodes in red are the most likely superspreaders in the spreading clusters in the epidemic network. The red arrows indicate the movement of superspreaders detected by DeepTrace as contact tracing continues.
@@ VI Conclusion
Digital contact tracing remains a crucial challenge in the pandemic response. This paper formulates both forward and backward contact tracing strategies as maximum likelihood estimations of the epidemic source, which are essential for identifying superspreaders. Starting from an index case, contact tracers explore the epidemic network using breadth-first or depth-first search to compute the graph center of the contact tracing subgraph, leading to optimal likelihood estimates for specific topologies. We introduce Algorithm DeepTrace, which employs a graph neural network with semi-supervised learning, incorporating pre-training and fine-tuning phases for large graphs through message passing. This approach enhances large-scale graph inference in computational epidemiology and connects to online graph exploration, providing low-complexity greedy algorithms for contact tracing. We validate our methodology by modeling real COVID-19 superspreading events in Hong Kong and Taiwan, showing that our algorithm outperforms existing heuristics.
To enhance digital contact tracing and address challenges like Antimicrobial Resistance (AMR), future work will focus on advanced graph exploration and adaptive learning to handle new epidemic and AMR networks. We aim to develop efficient decomposition techniques for optimization, extend Algorithm DeepTrace to process real-time data streams, and improve scalability for larger networks. By integrating pathogen resistance data, our approach can optimize contact tracing and contribute to scalable, data-driven surveillance systems for pandemics and AMR mitigation.
VII Acknowledgements
C. W. Tan acknowledges helpful discussions on the subject with O. T. Ng and K. Marimuthu of the Singapore National Centre for Infectious Diseases.
@@ A-A Proof of Theorem 1
Let $\mathbb{G}_{N}$ be a $d$-regular tree and $\overline{v}\in\mathbb{G}_{N}$ satisfies (5), we assume that the maximum distance from $\overline{v}$ to all leaves nodes is $k+1$. Then by (5), there are only two possible distances from $\overline{v}$ to all leaves, which are $k$ and $k+1$. For simplicity, let $T_{d,k}$ denote the perfect $d$-regular tree with $k$ levels, and we assume that the root level is 00. We denote $t^{u}_{v}$ as the size of the subtree rooted at $v$ after removing the edge $\{(u,v)\}$ from $\mathbb{G}_{N}$. We have the following properties of $\mathbb{G}_{N}$:
@@ A-C Proof of Property 2 of $\mathbb{G}_{N}$ used in Theorem 1
Let $\mathbb{G}_{N}$ be a $d$-regular tree and $\overline{v}\in\mathbb{G}_{N}$ satisfies (5), we assume that the maximum distance from $\overline{v}$ to all leaves nodes is $k+1$. Let $\overline{v}$ be the root of $\mathbb{G}_{N}$, and $v\in\mathbb{G}_{N}$ be an arbitrary node. If $v$ is a leaf or the root of $\mathbb{G}_{N}$, then the subtree rooted at $v$ satisfies (5). If $v$ is not a leaf or the root of $\mathbb{G}_{N}$, then we prove this property by contradiction. Assume that the subtree rooted at $v$, denoted as $T_{v}^{\overline{v}}$ does not satisfy (5), then there are two leaves $u,w\in T_{v}^{\overline{v}}$, satisfy $|\textup{dist}(u,v)-\textup{dist}(w,v)|>1$. since $\mathbb{G}_{N}$ is a tree, we have $\textup{dist}(\overline{v},u)=\textup{dist}(\overline{v},v)+\textup{dist}(v,u)$ and $\textup{dist}(\overline{v},w)=\textup{dist}(\overline{v},v)+\textup{dist}(v,w)$ which implies $|\textup{dist}(\overline{v},u)-\textup{dist}(\overline{v},w)|>1$ and contradicts to our assumption.
@@ A-D Proof of Theorem 2
Note that if a node in $\mathbb{G}_{N}$ with weight strictly less than $n/2$, then the node is the unique ML estimator on contact tracing network $\mathbb{G}_{N}$ [16, 17, 15]. Hence we have $j_{1}+1=2d_{1}-1.$. Moreover, since $(v=v^{*}_{j_{2}},v^{*}_{j_{2}+1})\in S_{2}$, we have
$j_{2}+1=2d_{2}-1>2(j_{1}+1)-1=2j_{1}+1.$
@@ A-E Proof of Lemma 1
Since the DFS contact tracing starts from the index case $v^{1}$, we can assume that the index case $v^{1}$ is the root of the DFS tree. Now, we can distinguish parent-node and child-node in each pair $(v^{*}_{j},v^{*}_{j+1})\in S$. Due to the property of the DFS tracing, a node is fully visited when all its descendants are fully visited; we have the following observation: If $(v^{*}_{j}=u,v^{*}_{j+1}=v)\in S_{2}\cup S_{3}$ and $v$ is the parent of $u$, then there is no other $i>j$ such that $(v^{*}_{i}=v,v^{*}_{i+1}=u)\in S_{2}\cup S_{3}$. Moreover, we can observe that:
@@
Since the estimated superspreader changes from $u$ to $v$ only when $P(\mathbb{G}_{N}\mid v)>P(\mathbb{G}_{N}\mid u)$. We can conclude that both type two and type three transitions must be followed by a type one transition, which means the size of the type one transition is more than type two and type three. Hence, we have $|S_{1}|\geq|S_{2}|+|S_{3}|$. Moreover, from the above two properties, we have the fact that for each $(u,v)\in S_{3}$, there is a corresponding $(v,u)\in S_{2}$, i.e., $|S_{2}|\geq|S_{3}|$. Combining the results above, we have $|S_{1}|\geq|S_{2}|\geq|S_{3}|$. Since $|S_{1}|+|S_{2}|+|S_{3}|=N-1$ and $|S_{2}|\geq|S_{3}|$, we have $|S_{1}|\geq\frac{N-1}{2}$ and $|S_{3}|\leq\frac{N-1}{4}$.
@@ A-F Proof of Theorem 3
As shown in (2), solving the MLE problem needs the information of all the permitted permutations for all the nodes in $G_{n}$. Since a newly infected node can only be infected by one of its infected neighbors, $v_{i+1}$ must be the neighbor of $v_{i}$ in any permitted permutation $\sigma$, and thus calculating $p(\sigma\mid v)$ with (4) is a recursive process of aggregating the information of degree $d(v_{l})$ and $\Phi_{l}$ from the $l$th layer’s neighbors of node $v$ in $\sigma$, which can be expressed as
@@
Note that (13) demonstrates the process of iterative aggregation to calculate $\mathbb{P}(G_{n}\mid v)$ for node $v$. Since $\mathbb{G}$ is connected, the shortest distance between any two nodes satisfies $\textup{dist}(u,v)\leq\textup{diam}(G)$. Thus, after some $\textup{diam}(G)$ steps of aggregation and combination, each node can obtain the feature information of all nodes in the graph. Furthermore, it is evident that both $f(\cdot)$ and $g(\cdot)$ are continuous functions. Consequently, based on the universal approximation theorem, these functions can be approximated to any desired degree of precision using standard multilayer perceptrons with hidden layers and non-constant, monotonically increasing activation functions [44].